OpenStudy (anonymous):

What are the solutions to 6x2 = 18x? Answer choices: x = 6, x = 3 x = 6, x = −3 x = 0, x = −3 x = 0, x = 3

OpenStudy (anonymous):

Um hi @Liltico

OpenStudy (campbell_st):

rewrite the equation \[6x^2 - 18x = 0\] take out the common factors \[6x^2 - 18x =0...... 6x(x - 3) = 0 \] then x = 0 or x = 3

OpenStudy (anonymous):

Thank you for explaining and answering! @campbell_st

OpenStudy (campbell_st):

hope it helps

OpenStudy (anonymous):

sorry..